OpenStudy (anonymous):

Solve for x: (Log base 3 x)^2 -log base 3 (x)^4 + 3 = 0

OpenStudy (anonymous):

I need your help @satellite73

hartnn (hartnn):

first use the property of log , for the middle term, that \(\log m^n = n\log m\)

OpenStudy (anonymous):

Hi. This is how I tried to do it

OpenStudy (anonymous):

2(log base3 x) - log base 3 (x)^4 + log base 3 (27) = log base 3 (1)

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

nopes (log m)^n is NOT n log m log [(m)^n] is actually n log m

OpenStudy (anonymous):

You're right.

OpenStudy (anonymous):

(Log base 3 x)^2 - Log base 3 x^4 + 3 = 0 (Log base 3 x)^2 - 4Log base 3 x + 3 = 0 Let a=Log base 3 a^2 - 4a + 3 = 0 (a-1)(a-3)=0 a=1 therefore Log base 3 x = 1 ------------->>> x = 3 a=3 therefore Log base 3 x = 3 ------------->>> x = 27

hartnn (hartnn):

yup, thats correct! god work :)